Determine the values of sin(2a), cos(2a) and tan(2a) , given cos(a) = -4/5 and Pi over 2 ≤ 0 ≤ Pi
1
Expert's answer
2020-10-25T19:04:29-0400

"cos(a)=-\frac{4}{5}"


lets find sin(a):

"sin^2(a)+cos^2(a)=1\implies sin(a)=\sqrt{1-cos^2(a)}"

"\sqrt{1-(-\frac{4}{5})^2}=\sqrt{\frac{9}{25}}=\pm\frac{3}{5}"

"\frac{\pi}{2}\leq a \leq \pi" then "\sin(a)=\frac{3}{5}"


tan(a):

"tan(a)=\frac{sin(a)}{cos(a)}=\frac{\frac{3}{5}}{-\frac{4}{5}}=-\frac{3}{4}"


using double angle formulas lets find sin(2a), cos(2a), tan(2a):


"sin(2a)=2sin(a)cos(a)=2\cdot\frac{3}{5}(-\frac{4}{5})=-\frac{24}{25}"


"cos(2a)=cos^2(a)-sin^2(a)=(-\frac{4}{5})^2-(\frac{3}{5})^2=\frac{16}{25}-\frac{9}{25}=\frac{7}{25}"


"tan(2a)=\frac{sin(2a)}{cos(2a)}=\frac{-\frac{24}{25}}{\frac{7}{25}}=-\frac{24}{7}"